OpenStudy (anonymous):

How do you write the equation y=-3x+7 in standard form?

OpenStudy (anonymous):

standard form: ax+by = c

OpenStudy (anonymous):

try to put it in that form by adding 3x to both sides

OpenStudy (anonymous):

oh i forgot there was also (1,3)

OpenStudy (anonymous):

can you rewrite the question again ?

OpenStudy (anonymous):

y=3x+2;(1,3)

OpenStudy (anonymous):

idk if you're looking for a parallel line or perpendicular line to (1,3) points

OpenStudy (anonymous):

it's in standard form

OpenStudy (anonymous):

my bad I gtogot that part.

OpenStudy (anonymous):

forgot*

OpenStudy (anonymous):

okay...so which is it?

OpenStudy (anonymous):

I don't inow it said write the equation of the line in standard form.

OpenStudy (anonymous):

if it's parallel, the slope is the same, if it's perpendicular, the slope is inverse with opposite sign. use y-y1 = m(x-x1) and plug in the point (1,3) alone with the slope. m = slope then put it in to slope intercept form of y = mx+b then put that in into standard form of ax+by =c

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (anonymous):

no problem